OpenStudy (anonymous):

Find the quotient: 16a²-b²/4a-b

OpenStudy (anonymous):

\[ \frac{16a^2-b^2}{4a-b}=\frac{(4a-b)(4a+b)}{4a-b}=4a+b; \;b\neq 4a\]